# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ 6.3 Link prediction and node classification for hierarchical graph data
Euc 59.8$\pm$2.0 32.5$\pm$1.1 92.0$\pm$0.0 60.9$\pm$3.4 83.3$\pm$0.1 48.2$\pm$0.7 82.5$\pm$0.3 23.8$\pm$0.7
Hyp 63.5$\pm$0.6 45.5$\pm$3.3 94.5$\pm$0.0 70.2$\pm$0.1 87.5$\pm$0.1 68.5$\pm$0.3 87.6$\pm$0.2 22.0$\pm$1.5
Euc-MIXED 49.6$\pm$1.1 35.2$\pm$3.4 91.5$\pm$0.1 68.3$\pm$2.3 86.0$\pm$1.3 63.0$\pm$0.3 84.4$\pm$0.2 46.1$\pm$0.4
Hyp-MIXED 55.1$\pm$1.3 56.9$\pm$1.5 93.3$\pm$0.0 69.6$\pm$0.1 83.8$\pm$0.3 73.9$\pm$0.2 85.6$\pm$0.5 45.9$\pm$0.3
MLP 72.6$\pm$0.6 28.8$\pm$2.5 89.8$\pm$0.5 68.6$\pm$0.6 84.1$\pm$0.9 72.4$\pm$0.2 83.1$\pm$0.5 51.5$\pm$1.0
HNNs 75.1$\pm$0.3 41.0$\pm$1.8 90.8$\pm$0.2 80.5$\pm$0.5 94.9$\pm$0.1 69.8$\pm$0.4 89.0$\pm$0.1 54.6$\pm$0.4
GCN 64.7$\pm$0.5 69.7$\pm$0.4 89.3$\pm$0.4 81.4$\pm$0.6 91.1$\pm$0.5 78.1$\pm$0.2 90.4$\pm$0.2 81.3$\pm$0.3
GAT 69.8$\pm$0.3 70.4$\pm$0.4 90.5$\pm$0.3 81.5$\pm$0.3 91.2$\pm$0.1 79.0$\pm$0.3 93.7$\pm$0.1 83.0$\pm$0.7
GraphSAGE 65.9$\pm$0.3 69.1$\pm$0.6 90.4$\pm$0.5 82.1$\pm$0.5 86.2$\pm$1.0 77.4$\pm$2.2 85.5$\pm$0.6 77.9$\pm$2.4
SGC 65.1$\pm$0.2 69.5$\pm$0.2 89.8$\pm$0.3 80.6$\pm$0.1 94.1$\pm$0.0 78.9$\pm$0.0 91.5$\pm$0.1 81.0$\pm$$0.1$
HGCNs 90.8$\pm$0.3 74.5$\pm$0.9 96.4$\pm$0.1 90.6$\pm$0.2 96.3$\pm$0.0 80.3$\pm$0.3 92.9$\pm$0.1 79.9$\pm$0.2
H2H-GCN 97.0$\pm$0.3 88.6$\pm$1.7 96.4$\pm$0.1 89.3$\pm$0.5 96.9$\pm$0.0 79.9$\pm$0.5 95.0$\pm$0.0 82.8$\pm$0.4
HGCN-Alg. 1 93.2$\pm$0.6 87.9$\pm$0.7 93.7$\pm$0.2 89.9$\pm$0.4 94.1$\pm$0.7 81.7$\pm$0.2 93.1$\pm$0.1 82.9$\pm$0.3
HGCN-Alg. 2 98.4$\pm$0.4 89.4$\pm$0.3 97.2$\pm$0.1 92.1$\pm$0.3 97.2$\pm$0.2 83.6$\pm$0.4 96.9$\pm$0.3 83.9$\pm$0.2
Finally, we show the utility of our methods as a pre-processing step for HGCNs [Chami et al., 2019, Dai et al., 2021], evaluated on LP and NC tasks. We adhere to the experimental setups and baselines used in these works to maintain consistency. For LP task, we use a Fermi-Dirac decoder [Krioukov et al., 2010, Nickel and Kiela, 2017] to compute probability scores for edges. Then, the networks are trained by minimizing cross-entropy loss with negative sampling. The performance of LP is assessed by measuring the area under the ROC curve (AUC). For NC task, we employ a centroid-based classification method [Liu et al., 2019], where softmax classifiers and cross-entropy loss functions are utilized. Additionally, an LP regularization objective is integrated into the NC task as in [Chami et al., 2019, Dai et al., 2021]. The NC task is evaluated using the F1 score for binary-class datasets and accuracy for multi-class datasets. We test four datasets [Sen et al., 2008, Yang et al., 2016, Chami et al., 2019], including CORA, PUBMED, DISEASE and AIRPOT. Descriptions of these datasets and their splits are included in App. E. We compare our methods with two shallow methods: Euclidean embedding (Euc) and Poincaré embedding (Hyp) [Nickel and Kiela, 2017]. We also include comparisons with the concatenation of shallow embeddings and node features, denoted as Euc-MIXED and Hyp-MIXED. Furthermore, we include multi-layer perceptron (MLP) and its hyperbolic extension, HNNs [Ganea et al., 2018], as well as four GNNs: GCN [Kipf and Welling, 2016], GAT [Veličković et al., 2017], GraphSAGE [Hamilton et al., 2017], and SGC [Wu et al., 2019]. Lastly, we include HGCNs [Chami et al., 2019] and H2H-GCN [Dai et al., 2021].
@@ 7 Conclusions
Limitation and future work. One limitation of our approach lies in the use of trees to represent data geometry. While this representation aligns with our assumption that the data exhibits underlying hierarchical structures, it may not generalize well to data supported on more complex geometries, e.g., spherical manifolds, spaces with mixed curvatures, or general graphs. In future work, we plan to explore more flexible geometric representation learning methods that can accommodate a broader class of data geometries, while utilizing the proposed iterative procedure between samples and features. We also plan to incorporate supervision or task-specific signals into the learning process. Finally, we plan to extend the approach to multi-mode data (e.g., tensor-valued inputs).
@@ Acknowledgments
We thank Gal Maman for proofreading an earlier version of the paper. The work of YEL and RT was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 802735-ERC-DIFFOP.
@@
  • Appendix C containes the proofs of Thm. 4.1 and Thm. 4.2, along with the supporting lemmas.
@@
  • Appendix F presents additional experimental results supporting our method. We first show a synthetic toy example simulating a video recommendation system for visualization. We then demonstrate the empirical uniqueness of the learned hierarchical representations under strong regularization. An ablation study evaluates the role of the iterative joint learning scheme in classification performance and contrasts it with applying wavelet filtering only after convergence. We further assess the effect of integrating wavelet filtering with alternative TWD variants, as well as with HGCNs without the iterative updates.
    We provide details on the sparse approximation analysis using Haar coefficients across iterations, as well as report the classification performance over the iterative learning scheme. Additionally, we include runtime analysis and explore the use of alternative regularizers within the iterative learning scheme. Finally, we present additional experiments on co-clustering tasks and incorporating fixed hierarchical distances in the iterative learning scheme for HGCNs.
  • Appendix G provides additional notes on the motivation for incorporating wavelet filters into our iterative learning scheme. We explain why wavelet filters are considered over Laplacian-based filtering in our context. Finally, we clarify the objective of our method and highlight the key differences between our approach and unsupervised ground metric learning with trees (Tree-WSV).
@@ A.1 Graph and tree
Graph and shortest path metric. A graph $G=(V,E,\mathbf{A})$ consists of a set of nodes $V$, edges $E$, and a edge weight matrix $\mathbf{A}\in\mathbb{R}^{m\times m}$. For any two nodes $j,j^{\prime}\in V$, the shortest path metric $d_{G}(j,j^{\prime})$ represents the minimum sum of edge weights along the path connecting node $j$ and node $j^{\prime}$.
Tree metric. A metric $d:V\times V\rightarrow\mathbb{R}$ is a tree metric if there exists a tree $T=(V,E,\mathbf{A})$ such that $d(j,j^{\prime})$ corresponds to the shortest path metric $d_{T}(j,j^{\prime})$ on $T$. Notably, tree metrics exhibit 00-hyperbolicity [Gromov, 1987], making them well-suited for representing hierarchical data.
Binary trees. Binary trees are a specific class of trees where each internal node has exactly two child nodes. In a balanced binary tree, internal nodes (except the root) have a degree of 3, while leaf nodes have a degree of 1. A rooted and balanced binary tree is a further refinement where one internal node, designated as the root, has a degree of 2, and all other internal nodes maintain a degree of 3. Note that a flexible tree can be represented as a binary tree through a transformation, e.g., left-child right-sibling [Cormen et al., 2022].
@@
Any tree metric (a metric derived from a tree graph where the distance between two nodes is the length of the unique path connecting them) can be isometrically embedded into an $\ell_{1}$ space [Bridson and Haefliger, 2013].
@@ A.2 Hyperbolic geometry
A hyperbolic geometry is a non-Euclidean geometry with a constant negative curvature. It is widely used in modeling structures with hierarchical or tree-like relationships [Ratcliffe et al., 1994]. There are four models commonly used to describe hyperbolic spaces. The Poincaré disk model maps the entire hyperbolic plane inside a unit disk, with geodesics represented as arcs orthogonal to the boundary or straight lines through the center. The Poincaré half-plane model uses the upper half of the Euclidean plane, where geodesics are semicircles orthogonal to the horizontal axis or vertical lines. The Klein model represents hyperbolic space in a unit disk but sacrifices angular accuracy, making it more suitable for some specific calculations. The hyperboloid model relies on a higher-dimensional Lorentzian inner product, offering computational efficiency.
In this work, we consider two equivalent models of hyperbolic space [Ratcliffe et al., 1994]: the Poincaré half-space and the hyperboloid model. The Poincaré half-space was used for embedding that reveals the hierarchical structure underlying high-dimensional data [Lin et al., 2023], while the hyperboloid model is advantageous for its simple closed-form Riemannian operations [Nickel and Kiela, 2018, Lin et al., 2021] used in hyperbolic graph convolutional networks [Chami et al., 2019, Dai et al., 2021].
@@
with the Riemannian metric tensor $ds^{2}=(d\mathbf{a}^{2}(1)+\ldots+d\mathbf{a}^{2}(d))/\mathbf{a}^{2}(d)$ [Beardon, 2012]. Let $\mathbb{L}^{d}$ denote the hyperboloid manifold in $d$ dimensions, defined by
@@
$\mathrm{Exp}_{\mathbf{b}_{1}}(\mathbf{v})=\cosh(||\mathbf{v}||_{\mathcal{L}})\mathbf{b}_{1}+\sinh(||\mathbf{v}||_{\mathcal{L}})\mathbf{v}/||\mathbf{v}||_{\mathcal{L}},$ (13)
@@ A.3 Diffusion geometry
Diffusion geometry [Coifman and Lafon, 2006] is a mathematical framework that analyzes high-dimensional data by capturing intrinsic geometric structures (i.e., manifolds with non-negative curvature) through diffusion processes. It is rooted in the study of diffusion propagation on graphs, manifolds, or general data spaces, where the spread of information or heat over time reflects the underlying connectivity and geometry. Diffusion geometry represents data as nodes in a graph and models their relationships using a diffusion operator. By simulating diffusion processes over this structure, the method identifies meaningful relationships based on proximity and connectivity in the diffusion space. We introduce the construction of the diffusion operator and its desired property below.
@@
The diffusion operator [Coifman and Lafon, 2006] is then constructed by $\mathbf{P}=\mathbf{K}\mathbf{D}^{-1}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $\mathbf{D}(j,j)=\sum_{l}\mathbf{K}(j,l)$. We remark that a density normalization affinity matrix can be considered to mitigate the effects of non-uniform data sampling [Coifman and Lafon, 2006]. Note that the diffusion operator $\mathbf{P}$ is column-stochastic, allowing it to be used as a transition probability matrix of a Markov chain on the graph $G$. Specifically, the vector $\mathbf{p}^{t}_{j}=\mathbf{P}^{t}\bm{\delta}_{j}$ is the propagated density after diffusion time $t\in\mathbb{R}_{+}$ of a density $\bm{\delta}_{j}$ initially concentrated at point $j$.
The diffusion operator has been demonstrated to have favorable convergence [Coifman and Lafon, 2006]. As $m\to\infty$ and $\epsilon\to 0$, the operator $\mathbf{P}^{t/\epsilon}$ converges pointwise to the Neumann heat kernel associated with the underlying manifold at the limits. This convergence indicates that the diffusion operator can be viewed as a discrete approximation of the continuous heat kernel, thereby effectively capturing the geometric structure of the underlying manifold in a finite-dimensional setting [Coifman and Lafon, 2006, Singer, 2006, Belkin and Niyogi, 2008].
@@ A.4 Tree construction based on hyperbolic and diffusion geometries
Recently, the work [Lin et al., 2023] introduced a tree construction method that recovers the latent hierarchical structure underlying high-dimensional data $\mathcal{Z}$ based on hyperbolic embeddings and diffusion geometry. Specifically, this latent hierarchical structure can be modeled by a weighted tree $T$, which can be viewed as a discretization of an underlying manifold $\mathcal{J}$. This manifold $\mathcal{J}$ is assumed to be a complete, simply connected Riemannian manifold with negative curvature, embedded in a high-dimensional ambient space $\mathbb{R}^{n}$, and equipped with a geodesic distance $d_{\mathcal{J}}$.
Consider the diffusion operator constructed as in App. A.3, in [Lin et al., 2023], the multi-scale diffusion densities $\bm{\mu}_{j}^{k}=\mathbf{P}^{2^{-k}}\bm{\delta}_{j}$ are considered with dyadic diffusion time steps $t=2^{-k}$ for $k\in\mathbb{Z}_{0}^{+}$, and are embedded into a Poincaré half-space by
@@
where ${\bm{\mu}_{j}^{k}}^{\circ 1/2}$ is the Hadamard root of $\bm{\mu}^{k}_{j}$. Note that diffusion geometry is effective in recovering the underlying manifold [Coifman and Lafon, 2006]; however, the study in [Lin et al., 2023] demonstrated that considering propagated densities with diffusion times on dyadic grids can capture the hidden hierarchical structures by incorporating local information from exponentially expanding neighborhoods around each point. The multi-scale hyperbolic embedding is a function $\mathtt{Embedding}:\mathcal{Z}\rightarrow\mathcal{M}$ defined by
@@
$d_{\mathcal{M}}(j,j^{\prime})=\sum_{k=0}^{K}~{}2\sinh^{-1}\left(2^{-k/2+1}\left\lVert\mathbf{y}^{k}_{j}-\mathbf{y}^{k}_{j^{\prime}}\right\rVert_{2}\right).$ (19)
Theorem A.1 (Theorem 1 in [Lin et al., 2023]).
@@
In [Lin et al., 2025], the multi-scale hyperbolic embedding is used to construct a binary tree $\mathcal{T}(\mathbf{M})$ with $m$ leaves, where each leaf corresponds to a data point in $\mathcal{Z}$. Pairs of points are merged based on the Riemannian mean [Fréchet, 1948] of the radius of the geodesic connecting the hyperbolic embeddings $\mathbf{y}_{j}^{k}$ and $\mathbf{y}_{j^{\prime}}^{k}$ for $k=\{0,\ldots,K\}$. Specifically, at the $k$-th hyperbolic embedding, the orthogonal projection on the $m+1$-axis in $\mathbb{R}^{m+1}$, which is the radius of the geodesic of $\mathbf{y}_{j}^{k}$ and $\mathbf{y}_{j^{\prime}}^{k}$, is given by
@@
Algorithm 3 Tree Construction Using Hyperbolic and Diffusion Geometry [Lin et al., 2025]
@@
   $\mathbf{K}\leftarrow\exp(-\mathbf{M}^{\circ 2}/\epsilon)$, $\mathbf{D}\leftarrow\mathtt{diag}(\mathbf{K})$, $\mathbf{P}\leftarrow\mathbf{KD}^{-1}$
@@
    $a_{j,j^{\prime}}=\sqrt[(K+1)]{\mathtt{proj}(\mathbf{y}_{j}^{0}\vee\mathbf{y}_{j^{\prime}}^{0})\cdots\mathtt{proj}(\mathbf{y}_{j}^{K}\vee\mathbf{y}_{j^{\prime}}^{K})}$
   $\mathcal{S}=\{(j,j^{\prime})|\text{sorted by }a_{j,j^{\prime}}\}$

### v3

@@ 6.3 Link prediction and node classification for hierarchical graph data
[…]
Finally, we show the utility of our methods as a pre-processing step for HGCNs [21, 45], evaluated on LP and NC tasks. We adhere to the experimental setups and baselines used in these works to maintain consistency. For LP task, we use a Fermi-Dirac decoder [93, 13] to compute probability scores for edges. Then, the networks are trained by minimizing cross-entropy loss with negative sampling. The performance of LP is assessed by measuring the area under the ROC curve (AUC). For NC task, we employ a centroid-based classification method [94], where softmax classifiers and cross-entropy loss functions are utilized. Additionally, an LP regularization objective is integrated into the NC task [21, 45]. The NC task is evaluated using the F1 score for binary-class datasets and accuracy for multi-class datasets. We test four datasets [9, 95, 21], including CORA, PUBMED, DISEASE and AIRPOT. Descriptions of these datasets and their splits are included in App. E. We compare our methods with two shallow methods: Euclidean embedding (Euc) and Poincaré embedding (Hyp) [13]. We also include comparisons with the concatenation of shallow embeddings and node features, denoted as Euc-MIXED and Hyp-MIXED. Furthermore, we include multi-layer perceptron (MLP) and its hyperbolic extension, HNNs [79], as well as four GNNs: GCN [96], GAT [97], GraphSAGE [98], and SGC [99]. Lastly, we include HGCNs [21] and H2H-GCN [45].
@@ 7 Conclusions
Limitations and future work.
One limitation of our approach lies in the use of trees to represent data geometry. While this representation aligns with our assumption that the data exhibits underlying hierarchical structures, it may not generalize well to data supported on more complex geometries, e.g., spherical manifolds [100], spaces with mixed curvatures [101], asymmetric data [102, 103, 104], or general graphs [105]. In future work, we plan to explore more flexible geometric representation methods that can accommodate a broader class of data geometries, while utilizing the proposed iterative procedure between samples and features. We also plan to explore a differentiable variant of TWD, such as the soft TWD [106], which could enable integration of our iterative process into neural architectures. We will further incorporate supervision or task-specific signals into the learning process. Finally, we plan to extend the approach to multi-way data (e.g., tensor-valued inputs).
@@ Acknowledgments
We thank the anonymous reviewers for their insightful feedback. We thank Gal Maman for proofreading this manuscript. The work of YEL and RT was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 802735-ERC-DIFFOP. The work of GM was supported by NSF CCF-2217058 and CCF-2403452. The work of RC was supported by AFOSR MURI 052721 and DARPA-PA-23-03.
@@
  • Appendix C contains the proofs of Thm. 1 and Thm. 2, along with the supporting lemmas.
@@
  • Appendix F presents additional experimental results supporting our method. We first show a synthetic toy example simulating a video recommendation system for visualization. We then demonstrate the empirical uniqueness of the learned hierarchical representations under strong regularization. An ablation study evaluates the role of the iterative joint learning scheme in classification performance and contrasts it with applying wavelet filtering only after convergence. We further assess the effect of integrating wavelet filtering with alternative TWD variants, as well as with HGCNs without the iterative updates. We provide details on the sparse approximation analysis using Haar coefficients across iterations, as well as report the classification performance over the iterative learning scheme. Additionally, we include runtime analysis and explore the use of alternative regularizers within the iterative learning scheme. We include additional experiments on co-clustering tasks and incorporating fixed hierarchical distances in the iterative learning scheme for HGCNs. We evaluate the effectiveness of our method as a preprocessing step across different neural network backbones. We further include a discussion and comparison with other tree distance metrics. Finally, we present the visualization of different iterations of the learned trees and how they evolve in our method on the toy example.
  • Appendix G provides additional notes on the motivation for incorporating wavelet filters into our iterative learning scheme. We explain why wavelet filters are considered over Laplacian-based filtering in our context. We clarify the objective of our method and highlight the key differences between our approach and unsupervised ground metric learning with trees (Tree-WSV). We discuss the interpretability of TWDs.
@@ A.1 Graph and tree
Graph and shortest path metric.
A graph $G=(V,E,\mathbf{A})$ consists of a set of nodes $V$, edges $E$, and a edge weight matrix $\mathbf{A}\in\mathbb{R}^{m\times m}$. For any two nodes $j,j^{\prime}\in V$, the shortest path metric $d_{G}(j,j^{\prime})$ represents the minimum sum of edge weights along the path connecting node $j$ and node $j^{\prime}$.
Tree metric.
A metric $d:V\times V\rightarrow\mathbb{R}$ is a tree metric if there exists a tree $T=(V,E,\mathbf{A})$ such that $d(j,j^{\prime})$ corresponds to the shortest path metric $d_{T}(j,j^{\prime})$ on $T$. Notably, tree metrics exhibit 0-hyperbolicity [107], making them well-suited for representing hierarchical data.
Binary trees.
Binary trees are a specific class of trees where each internal node has exactly two child nodes. In a balanced binary tree, internal nodes (except the root) have a degree of 3, while leaf nodes have a degree of 1. A rooted and balanced binary tree is a further refinement where one internal node, designated as the root, has a degree of 2, and all other internal nodes maintain a degree of 3. Note that a flexible tree can be represented as a binary tree through a transformation, e.g., left-child right-sibling [108].
@@
Any tree metric (a metric derived from a tree graph where the distance between two nodes is the length of the unique path connecting them) can be isometrically embedded into an $\ell_{1}$ space [109].
@@ A.2 Hyperbolic geometry
A hyperbolic geometry is a non-Euclidean geometry with a constant negative curvature. It is widely used in modeling structures with hierarchical or tree-like relationships [110]. There are four models commonly used to describe hyperbolic spaces. The Poincaré disk model maps the entire hyperbolic plane inside a unit disk, with geodesics represented as arcs orthogonal to the boundary or straight lines through the center. The Poincaré half-plane model uses the upper half of the Euclidean plane, where geodesics are semicircles orthogonal to the horizontal axis or vertical lines. The Klein model represents hyperbolic space in a unit disk but sacrifices angular accuracy, making it more suitable for some specific calculations. The hyperboloid model relies on a higher-dimensional Lorentzian inner product, offering computational efficiency.
In this work, we consider two equivalent models of hyperbolic space [110]: the Poincaré half-space and the hyperboloid model. The Poincaré half-space was used for embedding that reveals the hierarchical structure underlying high-dimensional data [50], while the hyperboloid model is advantageous for its simple closed-form Riemannian operations [14, 111] used in hyperbolic graph convolutional networks [21, 45].
@@
with the Riemannian metric tensor $ds^{2}=(d\mathbf{a}^{2}(1)+\ldots+d\mathbf{a}^{2}(d))/\mathbf{a}^{2}(d)$ [112]. Let $\mathbb{L}^{d}$ denote the hyperboloid manifold in $d$ dimensions, defined by
@@
[…]
@@ A.3 Diffusion geometry
Diffusion geometry [51] is a mathematical framework that analyzes high-dimensional data by capturing intrinsic geometric structures (i.e., manifolds with non-negative curvature) through diffusion processes. It is rooted in the study of diffusion propagation on graphs, manifolds, or general data spaces, where the spread of information or heat over time reflects the underlying connectivity and geometry. Diffusion geometry represents data as nodes in a graph and models their relationships using a diffusion operator. By simulating diffusion processes over this structure, the method identifies meaningful relationships based on proximity and connectivity in the diffusion space. We introduce the construction of the diffusion operator and its desired property below.
@@
The diffusion operator [51] is then constructed by $\mathbf{P}=\mathbf{K}\mathbf{D}^{-1}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $\mathbf{D}(j,j)=\sum_{l}\mathbf{K}(j,l)$. We remark that a density normalization affinity matrix can be considered to mitigate the effects of non-uniform data sampling [51]. Note that the diffusion operator $\mathbf{P}$ is column-stochastic, allowing it to be used as a transition probability matrix of a Markov chain on the graph $G$. Specifically, the vector $\mathbf{p}^{t}_{j}=\mathbf{P}^{t}\bm{\delta}_{j}$ is the propagated density after diffusion time $t\in\mathbb{R}_{+}$ of a density $\bm{\delta}_{j}$ initially concentrated at point $j$.
The diffusion operator has been demonstrated to have favorable convergence [51]. As $m\to\infty$ and $\epsilon\to 0$, the operator $\mathbf{P}^{t/\epsilon}$ converges pointwise to the Neumann heat kernel associated with the underlying manifold at the limits. This convergence indicates that the diffusion operator can be viewed as a discrete approximation of the continuous heat kernel, thereby effectively capturing the geometric structure of the underlying manifold in a finite-dimensional setting [51, 113, 114].
@@ A.4 Tree construction based on hyperbolic and diffusion geometries
Recently, the work [50] introduced a tree construction method that recovers the latent hierarchical structure underlying high-dimensional data $\mathcal{Z}$ based on hyperbolic embeddings and diffusion geometry. Specifically, this latent hierarchical structure can be modeled by a weighted tree $T$, which can be viewed as a discretization of an underlying manifold $\mathcal{J}$. This manifold $\mathcal{J}$ is assumed to be a complete, simply connected Riemannian manifold with negative curvature, embedded in a high-dimensional ambient space $\mathbb{R}^{n}$, and equipped with a geodesic distance $d_{\mathcal{J}}$.
Consider the diffusion operator constructed as in App. A.3, the multi-scale diffusion densities $\bm{\mu}_{j}^{k}=\mathbf{P}^{2^{-k}}\bm{\delta}_{j}$ are considered with dyadic diffusion time steps $t=2^{-k}$ for $k\in\mathbb{Z}_{0}^{+}$ [50, 115], and are embedded into a Poincaré half-space by
@@
where ${\bm{\mu}_{j}^{k}}^{\circ 1/2}$ is the Hadamard root of $\bm{\mu}^{k}_{j}$. Note that diffusion geometry is effective in recovering the underlying manifold [51]; however, the study [50] demonstrated that considering propagated densities with diffusion times on dyadic grids can capture the hidden hierarchical structures by incorporating local information from exponentially expanding neighborhoods around each point. The multi-scale hyperbolic embedding is a function $\mathtt{Embedding}:\mathcal{Z}\rightarrow\mathcal{M}$ defined by
@@
$d_{\mathcal{M}}(j,j^{\prime})=\hskip-2.84526pt\sum_{k=0}^{K}~2\sinh^{-1}\hskip-2.84526pt\left(2^{-k/2+1}\left\lVert\mathbf{y}^{k}_{j}-\mathbf{y}^{k}_{j^{\prime}}\right\rVert_{2}\right).$ (19)
Theorem A.1 (Theorem 1 [50]).
@@
The multi-scale hyperbolic embedding is used to construct a binary tree $\mathcal{T}(\mathbf{M})$ with $m$ leaves [37], where each leaf corresponds to a data point in $\mathcal{Z}$. Pairs of points are merged based on the Riemannian mean [116] of the radius of the geodesic connecting the hyperbolic embeddings $\mathbf{y}_{j}^{k}$ and $\mathbf{y}_{j^{\prime}}^{k}$ for $k=\{0,\ldots,K\}$. Specifically, at the $k$-th hyperbolic embedding, the orthogonal projection on the $m+1$-axis in $\mathbb{R}^{m+1}$, which is the radius of the geodesic of $\mathbf{y}_{j}^{k}$ and $\mathbf{y}_{j^{\prime}}^{k}$, is given by
@@
Algorithm 3 Tree Construction Using Hyperbolic and Diffusion Geometry [37]
@@
$\mathbf{K}\leftarrow\exp(-\mathbf{M}^{\circ 2}/\epsilon)$, $\mathbf{D}\leftarrow\mathtt{diag}(\mathbf{K})$, $\mathbf{P}\leftarrow\mathbf{K}\mathbf{D}^{-1}$
@@
  $a_{j,j^{\prime}}\leftarrow\sqrt[(K+1)]{\mathtt{proj}(\mathbf{y}_{j}^{0}\vee\mathbf{y}_{j^{\prime}}^{0})\cdots\mathtt{proj}(\mathbf{y}_{j}^{K}\vee\mathbf{y}_{j^{\prime}}^{K})}$
$\mathcal{S}\leftarrow\{(j,j^{\prime})\mid\text{sorted by }a_{j,j^{\prime}}\}$